A certain substance has a heat of vaporization of 41.25 kJ/mol.41.25 kJ/mol. At what Kelvin temperature will the vapor pressure be 5.505.50times higher than it was at 309 K?
Answer #1

   T1 = 309 K

     P2 = 5.50 x atm

     P1 = x

Hvap= 41.25 kJ/mol

R = 8.314 x 10^-3 kJ/K mol

from clausius -clayperon equation

ln(P2/P1) = ( Hvap/ R ) [1/T1 - 1/T2]

ln (5.50 / 1) = 41.25 / 8.314 x 10^-3 [1 / 309 - 1/ T2]

T2 = 345.7 K

temperature = 346 K